Two point charges Q1 = -25 micro coulombs and Q2 = +50 micro coulombs are separated by a distance of 12 cm. The electric field at the point P is zero (and is to the left of Q1). How far from (Q1) is P?


1
Expert's answer
2021-05-30T13:22:05-0400


Write the fields from both charges:


"E_1=\frac{kq_1}{r^2},\\\space\\
E_2=\frac{kq_2}{(x+r)^2},\\\space\\
E_1+E_2=0,\\\space\\
\frac{kq_1}{r^2}+\frac{kq_2}{(x+r)^2}=0,\\\space\\
q_1x^2+2q_1rx+(q_1+q_2)r^2=0,\\\space\\
r_{1,2}=\frac{-q_1x\pm\sqrt{(q_1x)^2-q_1(q_1+q_2)x^2}}{q_1+q_2}=\pm0.17\text{ m}."